An engine is tested by means of a water brake at 1000 rpm. The measured torque of the engine is 10000 Nm and the water consumption of the brake is 0.5 m3 /s, its inlet temperature being 20°C. 

Calculate the water temperature at exit, assuming that the whole of the engine power is ultimately  transformed into heat which is absorbed by the cooling water.


1
Expert's answer
2021-01-10T18:27:38-0500

"n=1000~rpm=16.67~rps,"

"M=10000~Nm,"

"K=\frac VT=0.5~\frac{m^3}{s},"

"t_1=20\u00b0C,"

"\rho=1000~\frac{kg}{m^3},"

"c=4200~\frac{J}{kg\cdot°C}."


"Q=cm(t_2-t_1)=PT,"

"P=M\omega,"

"cm(t_2-t_1)=M\omega T,"

"m=\rho V,~\omega=2\pi n,"

"t_2-t_1=\frac{2\pi nMT}{c\rho V}=\frac{2\pi nM}{c\rho K},"

"t_2=t_1+\frac{2\pi nM}{c\rho K}=20+\frac{2\cdot 3.14\cdot 16.67\cdot 10000}{4200\cdot 1000\cdot 0.5}=20.5~°C."